OpenStudy (anonymous):

what is the x intercept of the graph: f(x) = 3log(x + 5) + 2

OpenStudy (anonymous):

3log(x + 5) + 2 = 0 log(x + 5)^3 = 2 log(x + 5)^3 =2 log(10) log(x + 5)^3 =log(100) (x + 5)^3 =(100) finish it offf... I got to go.

OpenStudy (anonymous):

minor error at start 0 = 3 log(x+5) +2 3 log(x+5) = - 2 minus sign log (x+5) = -2/3 (x+5) = 10^(-2/3) x = 10^(-2/3) -5 etc.